# Designing Elegant Bell Inequalities

### v2

@@ Acknowledgements.
This research was supported by Korea Institute of Science and Technology Information(KISTI).(No. (KISTI)K24L4M1C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No. CAP22053-000). J. R. acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890).
Appendix A Symmetries on the coefficient matrices and optimal measurements
@@
We present the coefficient matrix for $S_{3}$ and MUBs and SICs that gives the optimal violation. Although the symmetries in $S_{3}$ can be expressed with an analytic equation in the main text, the explicit form of the coefficient matrix, MUBs and SICs would show the symmetries in the forms that can be seen at a glance. The coefficient matrices are given as $8\times 9$ matrices satisfying real-valuedness condition, $F_{2}=F_{1}^{*}$ as follows
@@
when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. One can find the first column generates the others following, $f_{dp+q,dr+s}^{1}=\omega^{-(ps+qr)}f_{3p+q,0}^{1}$.
@@
where each column of $\mbox{SIC}_{1}$ and $\mbox{SIC}_{2}$ defines SICs generated from $\left|\varphi_{1}\right>$ and $\left|\varphi_{2}\right>$ respectively. On the other side, the eigenstates of Alice’s optimal observables $A_{1}=Z$, $A_{3}=X$, $A_{4}=XZ$, $A_{5}=XZ^{2}$ respectively define the complete mutually unbiased bases,
@@
We note that the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs. It can be easily shown from the fact that they are unitary and up to phase equivalent to $A_{1}^{2},A_{3}^{2},A_{5}^{2},A_{4}^{2}$ respectively.
Appendix B Derivation of the quantum bound
Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is presented in this section. As shown in the main text, the SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as
by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$. The second term of the RHS of (13) is reduced to $d^{2}(d-1)/2\mathds{1}$ from the unitarity of $B_{y}^{n}$. The first term of the RHS is derived as the desirable form by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
$\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}$ (14)
where the Frobeniuns norm $\lVert F_{n}\rVert:=(\sum_{x,y}|f_{x,y}^{n}|^{2})^{1/2}$. Consequently, the upper bound in (13) is reduced to
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right]$ (15)
under conditions $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all $n,a\neq x$.

### v3

@@ 6 Conclusion
This research was supported by Korea Institute of Science and Technology Information(KISTI).(No. (KISTI)K25L5M2C2) and the National Research Council of Science & Technology(NST) grant by the Korea government (MSIT) (No. CAP22053-000). J. R. acknowledges support from the National Research Foundation (NRF) of Korea grant funded by the Korea Government (Grant No. NRF-2022M3K2A1083890).
Appendix A MUBs in Alice’s eigenspace
Complete MUBs are realized in the eigenspace of Alice’s $d^{2}-1$ measurements under condition (ii) of generalized EBI correlation. As previously noted in section 3, it is straightforwardly shown from $\mathcal{C}:=\{Z,X,XZ,\ldots,XZ^{d-1}\}\subset\mathcal{W}_{d}$ and the fact that the eigenstates of the $d+1$ unitary operators in $\mathcal{C}$ are complete MUBs for prime dimension [43]. We prove another notable property regarding MUBs on Alice’s eigenspace in the following proposition.
Proposition. The set of all eigenstates of Alice’s measurements defined in generalized EBI correlation are $(d-1)$-fold complete MUBs.
Proof. The measurements generating complete MUBs, $\mathcal{C}$, are divided into two subsets $\mathcal{C}_{1}:=\{Z\}$ and $\mathcal{C}_{2}:=\{XZ^{q}|q\in[0,d)\}$ such that $\mathcal{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$. The set of Alice’s measurements $\mathcal{A}:=\{X^{p}Z^{q}|(p,q)\neq(0,0)\}$ can be similarly divided with $\mathcal{A}_{1}:=\{Z^{q}|q\neq 0\}$ and $\mathcal{A}_{2}:=\{X^{p}Z^{q}|p\neq 0\}$, so that $\mathcal{A}=\mathcal{A}_{1}\cup\mathcal{A}_{2}$.
Then the above proposition is proved by
$V(\mathcal{C}_{1})$ $=V(\mathcal{A}_{1}),\,\mbox{and}$ (18)
$V(\mathcal{C}_{2})$ $=V(\mathcal{A}_{2})$ (19)
where $V(O)$ denotes a set of unit eigenvectors of all elements in a given set, $O$. The Eq.(18) are easily proved, as taking the power of a unitary matrix preserves the eigenspace.
Applying the same property of unitary matrices to the left-hand side (LHS) of the Eq.(19) derives $V(\mathcal{C}_{2})=V(\{(XZ^{q})^{p}|p\neq 0\})$. It also holds that $(XZ^{q})^{p}=\omega^{n}X^{p}Z^{\dot{(pq)}}$ from the canonical commutation relation, $Z^{q}X^{p}=\omega^{pq}X^{p}Z^{q}$, where $n(p,q):=qp(p-1)/2$ is an integer and the overdot denotes the smallest non-negative $d$ modulo value. Then, one obtains $V(\mathcal{C}_{2})=V(\{X^{p}Z^{\dot{(pq)}}|p\neq 0\})$. Consequently, the Eq.(19) can be converted into
$V(\{X^{p}Z^{\dot{(pq)}}|p\neq 0\})=V(\mathcal{A}_{2})=V(\{X^{p}Z^{q}|p\neq 0\}).$ (20)
This equation can be proved with the below equivalence
$q_{1}\neq q_{2}\Leftrightarrow\dot{(pq_{1})}\neq\dot{(pq_{2})}$ (21)
for $p\in(0,d)$ and $q_{1},q_{2}\in[0,d)$. It means that all different values of $q$ ($q_{1}$ and $q_{2}$) implies different values of $\dot{(pq_{1})}$ and $\dot{(pq_{2})}$. The situation results in the one-to-one correspondence between set of $q$ and set of $\dot{(pq)}$ for any given non-zero $p$. This is sufficient to proves Eq.(20).
We prove the rightward relation ($\Rightarrow$) first. The least non-negative $d$-modulo value of any non-negative integer $x$ can be expressed as $\dot{x}=x-d\lfloor x\rfloor$ with the quotient denoted by $\lfloor x\rfloor$. Using this expression, the RHS of (21) can be modified as
$\frac{p(q_{1}-q_{2})}{d}\neq\left\lfloor\frac{pq_{2}}{d}\right\rfloor-\left\lfloor\frac{pq_{1}}{d}\right\rfloor.$ (22)
Now suppose that the LHS of (21), $q_{1}\neq q_{2}$ Without loss of generality one can consider the case $q_{1}>q_{2}$. Then, as $d$ is prime and $p$ and $q_{1}-q_{2}$ are all smaller than $d$, the LHS of (22) is always a fractional number. While the RHS clearly is an integer. It proves (22), thus rightward relation. The leftward relation ($\Leftarrow$) is easily proved from its contraposition, $q_{1}=q_{2}\Rightarrow\dot{(pq_{1})}=\dot{(pq_{2})}$. $\blacksquare$
We note that the above proposition implies that any one of the eigenstates of Alice’s measurements fall into one of the eigenstates of $Z,X,XZ,\ldots,XZ^{d-1}$ that comprise the complete MUBs.
Appendix B Quantum bound
Detailed version of the derivation of quantum upper bound $\tilde{Q}$ is given in this section. As shown in the main text, the SOS decomposition ¯$\bar{\mathcal{B}}=\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified to
by substituting $P_{y}^{n}:=D_{y}^{n}\otimes\mathds{1}-\mathds{1}\otimes(B_{y}^{n})^{\dagger}$. The second term of the RHS of (23) is reduced to $d^{2}(d-1)/2\mathds{I}$ from the unitarity of $B_{y}^{n}$. The first term of the RHS is converted into an identity matrix multiplied by some real constant by substituting $D_{y}^{n}:=\sum_{x}f_{x,y}^{n}A_{x}^{n}$ into $\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$, as follows,
$\sum_{n,y}(D_{y}^{n})^{\dagger}D_{y}^{n}$ $=\sum_{n,y}\left(\sum_{a}f_{a,y}^{n}A_{a}^{n}\right)^{\dagger}\sum_{x}f_{x,y}^{n}A_{x}^{n}$
$=\sum_{n}\lVert F_{n}\rVert^{2}\mathds{1}+\sum_{n}\sum_{a\neq x}(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}(A_{a}^{n})^{\dagger}A_{x}^{n}.$ (24)
[…]
$\tilde{Q}:=\frac{1}{2}\left[\sum_{n}\lVert F_{n}\rVert^{2}+d^{2}(d-1)\right].$ (25)
Appendix C Coefficient matrix, MUBs and SICs in three dimension
@@
We present the coefficient matrix for $S_{3}$, and MUBs and SICs that gives the optimal violation. Although the symmetries in $S_{3}$ can be expressed with the analytic equation in the main text, the explicit forms of the coefficient matrix, MUBs and SICs would reveal the symmetries such that it can be seen at a glance. The coefficient matrices are given as $8\times 9$ matrices as follows:
@@
when $\lambda:=-i/\sqrt{3}$, $\mu:=\omega\lambda$ and $\nu:=\omega^{2}\lambda$. As expected, one can find that Eq.(12) holds. All the other coefficients is obtained from the realvaluedness condition, $F_{2}=F_{1}^{*}$.
@@
where the columns of $\mbox{SIC}_{\beta}$ correspond to the normalized vectors defining 9 SIC-POVM elements generated from $\left|\varphi_{\beta}\right>$. On the other side, the eigenstates of Alice’s optimal observables $A_{1}=Z$, $A_{3}=X$, $A_{4}=XZ$, $A_{5}=XZ^{2}$ respectively define the complete mutually unbiased bases,
@@
As previously discussed in Appendix A, the eigenvectors of other settings $A_{2},A_{6},A_{7},A_{8}$ also define the above four MUBs.